If the lengths of an isoselese are 6, 6 and 10 what is the angle of the obtuse angle?
Answer by Alan3354(69443)

If the lengths of an isoselese are 6, 6 and 10 what is the angle of the obtuse angle?
------------
The 2 acute angles are arccos(5/6) =~ 33.557 degs
Tha obtuse angle = 180 - 2*33.57 = ~112.885 degs